Jay-Zee Company makes an in-car navigation system. Next year, Jay-Zee plans to sell 20,000 units at a price of $360 each. Product costs include:

Direct materials $76.00
Direct labor $43.00
Variable overhead $11.00
Total fixed factory overhead $698,400

Variable selling expense is a commission of 5 percent of price; fixed selling and administrative expenses total $117,800.

Required:

1. Calculate the sales commission per unit sold. If required, round your answers to the nearest dollar. Use rounded answers in subsequent computations.
$ per unit

Calculate the contribution margin per unit.
$ per unit

2. How many units must Jay-Zee Company sell to break even? Round your answer to the nearest whole number.
units

Prepare an income statement for the calculated number of units. If an amount is zero, enter "0". Do NOT round Break-even units and, if required, round your answer to the nearest dollar.

Jay-Zee Company
Income Statement
$
$
$

3. Calculate the number of units Jay-Zee Company must sell to achieve target operating income (profit) of $384,144. Round your answer to the nearest whole number.
units

4. What if the Jay-Zee Company wanted to achieve a target operating income of $371,000? Would the number of units needed increase or decrease compared to your answer in Requirement 3? Round your answer to the nearest whole number.

Compute the number of units needed for the new target operating income.
units
Answer #1

1) Sales comission per unit = 360*5% = 18 per unit

Contribution margin per unit = 360-76-43-11-18 = 212

2) Break even point = (698400+117800)/212 = 3850 Units

Prepare an income statement for the calculated number of units. If an amount is zero, enter "0". Do NOT round Break-even units and, if required, round your answer to the nearest dollar.

Jay-Zee Company
Income Statement
Sales $1386000
Variable cost 569800
Contribution margin $816200
Fixed cost 816200
Income (loss) $0

3) Target unnit = (816200+384144)/212 = 5662 Units

4) No of unit will decrease

No of units = (816200+371000)/212= 5600 Units